Solution
= Solution
For independent fair dice, the <probability> of avoiding one on each die is $5/6$. Therefore
$$
\boxed{P(\text{neither is one})=\left(\frac56\right)^2=\frac{25}{36}}.
$$
The <complement of an event> gives
$$
\boxed{P(\text{at least one is one})=1-\frac{25}{36}=\frac{11}{36}}.
$$